# SMARTe-VR: Student Monitoring and Adaptive Response Technology for e-learning in Virtual Reality

### v1

@@ Abstract
This work introduces SMARTe-VR, a platform for student monitoring in an immersive virtual reality environment designed for online education. SMARTe-VR is aimed to gather data for adaptive learning, focusing on facial biometrics and learning metadata. The platform allows instructors to create tailored learning sessions with video lectures, featuring an interface with an Auto QA system to evaluate understanding, interaction tools (e.g., textbook highlighting and lecture tagging), and real-time feedback. Additionally, we release a dataset containing 5 research challenges with data from 10 users in VR-based TOEIC sessions. This dataset, spanning over 25 hours, includes facial features, learning metadata, 450 responses, question difficulty levels, concept tags, and understanding labels. Alongside the database, we present preliminary experiments using Item Response Theory models, adapted for understanding detection using facial features. Two architectures were explored: a Temporal Convolutional Network for local features and a Multilayer Perceptron for global features.
1 Introduction
@@
In recent years, education has undergone a transformation driven by the digital revolution. A key factor in this change has been the growing adoption of online education or e-learning, further accelerated by the COVID-19 pandemic, which forced a transition toward remote teaching models. This modality is expected to continue growing exponentially over the next 20 years (Chen 2018; Daza et al. 2023b).
@@
Online education has traditionally been delivered through devices such as PC, laptops, and even tablets. However, as emerging technologies like virtual reality (VR) have become more affordable in recent years, a question arises: could these environments be more engaging than traditional e-learning? Through the metaverse, VR offers a sense of presence and interactivity that has often been missing in online education (Nagao 2023).
@@
One of VR’s key strengths is its ability to more easily capture biometric and behavioral data, thanks to sensors placed on the face, hands, and other areas. This level of data collection is challenging to achieve in conventional e-learning environments. However, VR also presents challenges like fatigue and cybersickness after prolonged use, raising questions about its viability in education (Souchet et al. 2023).
@@
This data collection capability not only facilitates the development and adoption of new learning systems based on biometric signal analysis, behavioral data, and learning analytics, but also addresses specific challenges in online education. Among these challenges is the lack of direct contact between teachers and students, which complicates aspects such as personalized learning (Kebritchi, Lipschuetz, and Santiague 2017).
@@
VR has the potential to enhance educational quality through the implementation of Adaptive Learning (AL) systems (Nagao 2019; Barbosa et al. 2024). Learning analytics tools, such as Item Response Theory (IRT), play a crucial role in these systems, as they enable continuous and personalized assessment of students’ knowledge and performance (Zhou, Suzuki, and Kumano 2024).
@@
The emotional and cognitive state of a student directly impacts their understanding and academic performance (Joshi et al. 2019; Bala et al. 2023). For this reason, recent research has expanded the scope of Learning Analytics (LA) into Multimodal Modal Learning Analytics (MMLA). Supported by advanced technologies such as deep learning, these approaches have achieved state-of-the-art results in predicting academic performance or knowledge. In particular, facial biometric signals have demonstrated a clear relationship with factors like attention, stress, and emotions, highlighting their potential for optimizing adaptive and personalized learning (Hernandez-Ortega et al. 2020a; Daza et al. 2023b; Kovanovic et al. 2023). VR offers a significant advantage in capturing facial expressions more easily and in an environment perceived as more realistic by the user, through VR headsets with facial sensors that allow detailed and continuous tracking of user interaction.
@@
However, current public databases in e-learning have several limitations:
@@
  • Most databases are not multimodal; they do not include biometric information or behavioral data.
@@
These limitations hinder the study of AL systems in VR, particularly for data-driven multimodal models. They also prevent comprehensive comparisons of the advantages and drawbacks relative to traditional e-learning environments.
@@
The contributions of this work are as follows: i) We present a survey of state-of-the-art AL and public e-learning databases; ii) We developed a conference visualization platform for VR, integrating an Auto QA system to assess student comprehension. This platform enables continuous monitoring by capturing metadata and facial expressions; iii) We designed an online session protocol, which includes an initial video lecture paired with a pretest, 10 video lectures with accompanying questions of varying difficulty levels, along with an automated note system using ChatGPT. This system allows students to quickly highlight information in VR environments using a color-coding scheme, facilitating review; iv) We have released an initial subset of the database, featuring 10 users in VR. Each LS lasts 2.5 hours, capturing metadata and facial features. The database was carefully designed to create a balanced dataset, specifically intended to train and evaluate data-driven interpretability models for AL; and v) We conducted preliminary experiments based on IRT systems to validate the dataset, using both local and global facial features collected during the lecture videos to predict whether the student has understood the content.
@@
The remainder of the paper is organized as follows: Section 2 provides a summary of related work on AL with a focus on IRT and public databases in e-learning. Section 3 presents the VR conference visualization platform, the published database, and the proposed evaluation challenges for the community. Section 4 outlines the experiments and results. Finally, conclusions and directions for future work are discussed in Section 5.
2 Related Work
@@
AL is an educational approach that personalises the learning experience for each student by dynamically adjusting content, pacing, and learning pathways based on individual needs and responses (Barbosa et al. 2024). This is relevant in e-learning, where students often receive identical resources without considering prior knowledge or skills, and direct interaction is limited (McCusker et al. 2013). AL has gained momentum in recent years, driven by advances in Artificial Intelligence (AI), Machine Learning, and data analytics.
2.1 Approaches in Adaptive Learning
@@
Traditionally, learning style models, such as Felder-Silverman and Kolb (Felder 2002; Kolb et al. 2007), have been used to personalize teaching; however, their effectiveness has been questioned due to a lack of solid empirical evidence. These models classify students according to sensory and cognitive preferences (An and Carr 2017). In response, modern adaptive systems prioritize objective data analysis, such as performance analytics and biometrics.
Modern adaptive education systems aim to model an abstract representation of a student’s skill or knowledge level based on their performance and interactions with the educational system. This is a continuous process, as students evolve over time (Desmarais and Baker 2012). The technological foundation lies in LA, an area gathering and analysing data on student interactions and engagement, providing real-time insights for AL adjustments (Chatti et al. 2012). The field is advancing towards MMLA, integrating diverse data sources like biometrics, behaviors, and metadata to offer a more holistic view of the educational process (Hernandez-Ortega et al. 2020a; Daza et al. 2023b). Three student models have been widely developed:
  • Item Response Theory (IRT): A statistical model designed to evaluate both student ability ($\theta$) and question difficulty ($b$) based on their responses, making it highly suitable for creating adaptive tests. IRT models utilize estimation methods such as Maximum Likelihood or Bayesian approaches, often employing iterative techniques for parameter estimation. IRT models vary based on the parameters they incorporate. The most common version is the 1PL (Rasch model), which includes only the difficulty parameter ($b$) for each item.
    Since traditional IRT models typically assume that examinees’ abilities are randomly sampled from a standard distribution, recent research has explored combining IRT with deep learning techniques. A notable example is the Deep-IRT model (Tsutsumi, Kinoshita, and Ueno 2021), which introduces two twin neural networks: one for modeling student ability and another for item difficulty. The network inputs are one-hot vectors for the question and the student. This model outperforms traditional ones across diverse datasets, predicting unknown item responses more accurately from previous histories and proving effective even for small datasets. However, it may encounter challenges with sparse data.
    Recent studies have adapted the 1PL model with deep learning to incorporate biometric signals, such as facial expressions, for predicting an additional parameter: the student state. One prominent example is the SAD-IRT model (Zhou, Suzuki, and Kumano 2024), which adjusts item difficulty dynamically based on students’ cognitive and emotional states. Built upon Deep-IRT, SAD-IRT adds a branch with spatial (MARLIN) (Cai et al. 2023) and temporal feature (Temporal Convolutional Network, TCN) extractors. This branch processes facial images captured while students solve test items to derive the student state parameter. Evaluated on a dataset of 20 participants and 1,000 responses, SAD-IRT outperformed both the DEEP-IRT model and facial video-only models.
@@
    ASSISTments 46,674 No No Eval. only Math responses (1 year); thousands of questions
    Junyi 247,606 No No Eval. only Math responses; 722 questions
@@
    EdNet (KT4) 297,915 No No Logs & Eval. TOEIC exam responses with logs of consulted material (15 months); thousands of questions
@@
    SAD-IRT 20 No Yes Eval. only 17 hours of assessments with facial tracking; 1,000 multiple-choice responses
@@
    edBB 60 No Yes Eval. only 10 hrs assessment, focused on behavioral detection
@@
    Ours 10 Yes Yes Class & Eval. 25 hours in VR, covering learning and assessment with facial tracking; 450 multiple-choice responses
    Table 1: Summary of public datasets in education (ASSISTments Foundation 2013; Chang et al. 2015; Choi et al. 2020; Zhou, Suzuki, and Kumano 2024; Hernandez-Ortega et al. 2020a), and SMARTe-VR. Bio. $=$ Biometrics, Eval. $=$ Evaluation, LS $=$ Learning Session and Logs $=$ Recorded material consulted by students.
@@
  • Knowledge Tracing (KT): Models the evolution of a student’s knowledge by updating an internal ’knowledge state’ based on responses over a sequence of questions, using techniques like the Hidden Markov Model. Unlike IRT, KT focuses on sequential learning rather than static skills, without accounting for question difficulty or individual ability. Advanced KT models, such as Deep KT, leverage deep learning for enhanced response prediction accuracy, though they lack the interpretability of IRT (Zhou, Suzuki, and Kumano 2024). Combined IRT-KT methods have shown promising results, alongside recent efforts to integrate biometric and behavioral data to capture affective states (Ghosh, Heffernan, and Lan 2020).
@@
  • Biometric and Behaviour-Based Learning Recognition (BBLR): These emerging models assess the student’s level of understanding during the session, without requiring constant feedback through answers. They integrate biometric and behavioral data—such as facial expressions (Zhou, Suzuki, and Kumano 2024), eye blinks (Daza et al. 2020, 2024c), keystroke dynamics (Morales et al. 2015, 2016), heart rate (Hernandez-Ortega et al. 2020b), head pose (Daza et al. 2024a; Becerra et al. 2024) and more—to estimate cognitive load (Daza et al. 2022, 2023a, 2024b), emotions (Bala et al. 2023), etc., linking these states to learning levels. Unlike traditional approaches such as IRT and KT, these models enable real-time feedback, adjusting content and pacing according to the student’s state. Their main challenge lies in a lack of interpretability due to complex deep learning architectures, which do not always clearly interpret the factors contributing to academic performance.
2.2 Adaptive Learning Platforms
@@
Several educational platforms have adopted AL models. For example, Duolingo uses IRT models to optimize language learning (Maris 2020), and platforms like ASSISTments and Coursera employ KT models or similar methods (Feng, Heffernan, and Koedinger 2009). Additionally, public institutions have developed their own AL systems, including:
@@
Learning Evidence Analytics Framework (LEAF): Developed at Kyoto University (Ogata et al. 2022), LEAF applies a learning analytics approach to collect, analyse, and visualise educational data. It includes tools such as BookRoll, Log Pallet, and LAView for real-time data visualization and feedback. Through AI-driven big data analysis, LEAF provides personalised learning support, enabling immediate feedback for students and instructors.
@@
edBB: Developed at the Autonomous University of Madrid (Hernandez-Ortega et al. 2020a; Daza et al. 2023b), edBB focuses on learning recognition through biometric and behavioral analysis, using CNN and RNN models to estimate attention, heart rate, and more. Integrated with M2LADS (Becerra et al. 2023), it offers instructors a comprehensive view of student progress.
2.3 Public State of the Art Database

### v3

@@ Abstract.
This work introduces SMARTe-VR, a platform for student monitoring in an immersive virtual reality environment designed for online education. SMARTe-VR aims to collect data for adaptive learning, focusing on facial biometrics and learning metadata. The platform allows instructors to create customized learning sessions with video lectures, featuring an interface with an AutoQA system to evaluate understanding, interaction tools (for example, textbook highlighting and lecture tagging), and real-time feedback. Furthermore, we released a dataset that contains 5 research challenges with data from 10 users in VR-based TOEIC sessions. This data set, which spans more than 25 hours, includes facial features, learning metadata, 450 responses, difficulty levels of the questions, concept tags, and understanding labels. Alongside the database, we present preliminary experiments using Item Response Theory models, adapted for understanding detection using facial features. Two architectures were explored: a Temporal Convolutional Network for local features and a Multilayer Perceptron for global features.
E-learning Platform, Virtual Reality, Facial Biometrics and Behavior, Item Response Theory, Deep Learning
journalyear: 2025copyright: acmlicensedconference: Proceedings of the International Workshop on Intelligent Immersification in the Metaverse: AI-Driven Immersive Multimedia; October 27–28, 2025; Dublin, Irelandbooktitle: Proceedings of the International Workshop on Intelligent Immersification in the Metaverse: AI-Driven Immersive Multimedia (I2M-$M$ ’25), October 27–28, 2025, Dublin, Irelandccs: Human-centered computing Virtual realityccs: Applied computing E-learningccs: Computing methodologies Machine learning
Figure 1. Diagram of the SMARTe-VR platform for e-learning in immersive virtual reality. The figure illustrates the key features of SMARTe-VR, including its ability to capture multimodal data, such as facial expressions and learning metadata, during VR sessions to support adaptive learning. It also highlights the collected dataset, which is associated with five proposed challenges for the community, as well as the Deep IRT-based models developed to address Challenge 1: Understanding Estimation during lecture sessions using biometric data. The visual elements align with the main contributions of the paper (i–v).
[…]
Abstract diagram of the SMARTe-VR platform, including the database and Deep-IRT models used to estimate student understanding. The visual elements align with the main contributions of the paper (i–v).
1. Introduction
@@
In recent years, education has undergone a transformation driven by the digital revolution. A key factor in this change has been the growing adoption of online education or e-learning, further accelerated by the COVID-19 pandemic, which forced a transition to remote teaching models. This modality is expected to grow exponentially in the next 20 years (Chen, 2018; Daza et al., 2023b).
@@
Online education has traditionally been delivered via devices such as PC, laptop, and even tablets. However, as emerging technologies such as virtual reality (VR) have become more affordable in recent years, a question arises: Could these environments be more engaging than traditional e-learning? Through the metaverse, VR offers a sense of presence and interactivity that has often been missing in online education (Nagao, 2023).
@@
One of VR’s key strengths is its ability to more easily capture biometric and behavioral data, thanks to sensors placed on the face, hands, and other areas. This level of data collection is challenging to achieve in conventional e-learning environments. However, VR also presents challenges such as fatigue and cybersickness after prolonged use, raising questions about its viability in education  (Souchet et al., 2023).
@@
This data collection capability not only facilitates the development and adoption of new learning systems based on biometric signal analysis, behavioral data, and learning analytics, but also addresses specific challenges in online education. Among these challenges is the lack of direct contact between teachers and students, which complicates aspects such as personalized learning  (Kebritchi et al., 2017).
@@
VR has the potential to improve educational quality through the implementation of Adaptive Learning (AL) systems (Nagao, 2019; Barbosa et al., 2024). Learning analytics tools, such as Item Response Theory (IRT), play a crucial role in these systems, as they enable continuous and personalized assessment of students’ knowledge and performance (Zhou et al., 2024).
@@
The emotional and cognitive state influences their understanding and academic performance (Joshi et al., 2019; Bala et al., 2023). For this reason, recent research has expanded the scope of Learning Analytics (LA) to Multimodal Learning Analytics (MMLA). Powered by advanced technologies such as deep learning, these approaches have achieved state-of-the-art results in predicting academic performance or knowledge. In particular, facial biometric signals have shown a clear relationship with factors such as attention, stress, and emotions, highlighting their potential to optimize adaptive and personalized learning (Hernandez-Ortega et al., 2020a; Daza et al., 2023b; Kovanovic et al., 2023). VR offers a significant advantage in capturing facial expressions more easily and in an environment perceived as more realistic by the user, through VR headsets with facial sensors that allow detailed and continuous tracking of user interaction.
@@
However, current public databases in e-learning have several limitations.
@@
  • Most databases are not multimodal; they do not include biometric or behavioral information.
@@
These limitations hinder the study of adaptive learning (AL) systems in VR, particularly for data-driven multimodal models. They also prevent comprehensive comparisons of the advantages and drawbacks relative to traditional e-learning environments.
@@
The contributions of this work are shown in Fig. 1 and detailed as follows: i) We present a survey of state-of-the-art AL and public e-learning databases. ii) We develop a visualization platform for VR, integrating an AutoQA system to assess student comprehension. This platform enables continuous monitoring by capturing metadata and facial expressions. iii) We designed an online session protocol, which includes an initial video lecture paired with a pretest, 10 video lectures with accompanying questions of varying difficulty levels, along with an automated note system using ChatGPT. This system allows students to quickly highlight information in VR environments using a color-coding scheme, facilitating review. iv) We have released an initial subset of the database (SMARTe-VR-db [^1] Available at: https://github.com/BLANCELIN1/SMARTe-VR-DB), featuring 10 VR users. Each Learning Session (LS) lasts 2.5 hours and captures metadata and facial features. The database was carefully designed to create a balanced dataset, specifically designed to train and evaluate data-driven interpretability models for AL. And v)  we conducted preliminary experiments based on Item Response Theory (IRT) systems to validate the data set, using local and global facial features collected during lecture videos to predict whether the student has understood the content.
@@
The remainder of the paper is organized as follows. Section 2 provides a summary of related work on AL with a focus on IRT and public databases in e-learning. Section 3 presents the VR visualization platform, the published database, and the evaluation challenges proposed for the community. Section 4 outlines the experiments and results. Finally, conclusions and directions for future work are discussed in Section 5.
2. Related Work: Adaptive Learning
@@
Adaptive Learning (AL) is an educational approach that personalizes the learning experience for each student by dynamically adjusting content, pace, and learning pathways according to individual needs and responses (Barbosa et al., 2024). This is relevant in e-learning, where students often receive identical resources without considering prior knowledge or skills, and direct interaction is limited (McCusker et al., 2013). AL has gained momentum in recent years, driven by advances in Artificial Intelligence (AI), Machine Learning, and data analytics.
2.1. Adaptive Learning: Approaches
@@
Traditionally, learning style models, such as Felder-Silverman and Kolb (Felder and Silverman, 1988; Kolb et al., 2007), have been used to personalize teaching; however, their effectiveness has been questioned due to the lack of solid empirical evidence. These models classify students according to sensory and cognitive preferences (An and Carr, 2017). In response, modern adaptive systems prioritize objective data analysis, such as performance analytics and biometrics.
Table 1. Summary of existing datasets for research in adaptive learning (AL) (ASSISTments Foundation, 2013; Chang et al., 2015; Choi et al., 2020; Zhou et al., 2024; Hernandez-Ortega et al., 2020a; Daza et al., 2025a) in comparison to our contributed SMARTe-VR-db. Bio. == Biometrics, Eval. == Evaluation, LS == Learning Session, and Logs == Recorded material consulted by students.
@@
ASSISTments (ASSISTments Foundation, 2013) 46,674 No No Eval. only Math responses (1 year); thousands of questions
Junyi (Chang et al., 2015) 247,606 No No Eval. only Math responses; 722 questions
@@
EdNet (KT4) (Choi et al., 2020) 297,915 No No Logs & Eval. TOEIC exam responses with logs of consulted material (15 months); thousands of questions
@@
SAD-IRT (Zhou et al., 2024) 20 No Yes Eval. only 17 hours of assessments with facial tracking; 1,000 multiple-choice responses
@@
edBB (Hernandez-Ortega et al., 2020a) 60 No Yes Eval. only 10 hrs assessment, focused on behavioral detection
IMPROVE (Daza et al., 2025a) 120 No Yes Class & Eval. Primarily focused on effects of phone use on learning
@@
SMARTe-VR-db (Ours) 10 Yes Yes Class & Eval. 25 hours in VR, covering learning and assessment with facial tracking; 450 multiple-choice responses
Modern adaptive education systems aim to model an abstract representation of a student’s skill or level of knowledge based on their performance and interactions with the educational system. This is a continuous process, as the students evolve over time (Desmarais and Baker, 2012). The technological foundation lies in LA, an area that collects and analyzes data on student interactions and engagement, providing real-time insights for AL adjustments (Chatti et al., 2012). The field is advancing towards MMLA, integrating diverse data sources such as biometrics, behaviors, and metadata to offer a more holistic view of the educational process (Hernandez-Ortega et al., 2020a; Daza et al., 2023b). Three student models have been widely developed:
  • Item Response Theory (IRT): A statistical model designed to evaluate both student ability ($\theta$) and question difficulty ($b$) based on their responses, making it highly suitable for creating adaptive tests. IRT models utilize estimation methods such as Maximum Likelihood Estimation (MLE) or Bayesian inference, often employing iterative techniques for parameter estimation. IRT models vary according to the parameters they incorporate. The most common version is the 1PL (Rasch model), which includes only the difficulty parameter ($b$) for each item (Baker, 2001; Himelfarb, 2019).
    Since traditional IRT models typically assume that examinees’ abilities are randomly sampled from a standard distribution, recent research has explored combining IRT with deep learning techniques. A notable example is the Deep-IRT model (Tsutsumi et al., 2021), which introduces two twin neural networks: one to model the ability of the student and the other to model the difficulty of the elements. The network inputs are one-hot vectors for the question and the student. This model outperforms traditional models across diverse datasets, predicting unknown item responses more accurately from previous histories, and proving effective even for small datasets. However, it may encounter challenges with sparse data.
    Recent studies have adapted the 1PL model with deep learning to incorporate biometric signals, such as facial expressions, to predict an additional parameter: the student’s state. One prominent example is the SAD-IRT model (Zhou et al., 2024), which adjusts item difficulty dynamically based on students’ cognitive and emotional states. Based on Deep-IRT, SAD-IRT adds a branch with spatial extractors (MARLIN) (Cai et al., 2023) and temporal features (Temporal Convolutional Network, TCN). This branch processes facial images captured while students solve test items to derive the student state parameter. Evaluated on a dataset of 20 participants and 1,000 responses, SAD-IRT outperformed both the DEEP-IRT model and the facial video-only models.
@@
  • Knowledge Tracing (KT): Models the evolution of a student’s knowledge by updating an internal ”knowledge state” based on responses over a sequence of questions, using techniques like the Hidden Markov Model. Unlike IRT, KT focuses on sequential learning rather than static skills, without accounting for the difficulty of the question or individual ability. Advanced KT models, such as Deep KT, leverage deep learning for improved response prediction accuracy, although they lack the interpretability of IRT (Zhou et al., 2024). Combined IRT-KT methods have shown promising results, in addition to recent efforts to integrate biometric and behavioral data to capture affective states (Ghosh et al., 2020).
@@
  • Biometrics- and Behaviour-Based Learning Recognition (BBLR): These emerging models assess the student’s level of understanding during the session, without requiring constant feedback through answers. They integrate biometric and behavioral data, such as facial expressions (Zhou et al., 2024), eye blinks (Daza et al., 2020, 2024b), visual attention (Navarro et al., 2024), keystroke dynamics (Morales et al., 2015, 2016), heart rate (Hernandez-Ortega et al., 2020b), head pose (Daza et al., 2025a; Becerra et al., 2024), etc., to estimate cognitive load (Daza et al., 2022, 2023a, 2024a), emotions (Bala et al., 2023), etc., linking these states to learning levels. Unlike traditional approaches such as IRT and KT, these models enable real-time feedback, adjusting content, and pacing according to the student’s state. Their main challenge lies in a lack of interpretability due to complex deep learning architectures, which do not always clearly interpret the factors that contribute to academic performance.
2.2. Adaptive Learning: Platforms
@@
Several educational platforms have adopted AL models. For example, Duolingo uses IRT models to optimize language learning (Maris, 2020), and platforms like ASSISTments and Coursera employ KT models or similar methods (Feng et al., 2009). Some public institutions have also developed AL systems, e.g.:
@@
  • Learning Evidence Analytics Framework (LEAF): Developed at Kyoto University (Ogata et al., 2022), LEAF applies learning analytics to collect, analyze, and visualize educational data. Includes tools such as BookRoll, Log Pallet, and LAView for real-time data visualization and feedback. Through AI-driven big data analysis, LEAF provides personalized learning support, enabling immediate feedback for students and instructors.
@@
  • edBB: Developed at the Autonomous University of Madrid (Hernandez-Ortega et al., 2020a; Daza et al., 2023b), edBB focuses on learning recognition through biometric and behavioral analysis, using CNN and RNN models to estimate attention, heart rate, and more. Integrated with M2LADS (Becerra et al., 2023b, a, 2025a), it offers instructors a comprehensive view of student progress.
2.3. Existing Research Datasets